State the circle and radius of the circle given by (x-2)^2+(y+3)^2=20
Found 2 solutions by Nate, jake_6233:
Answer by Nate(3500)

Use standard form: (x - h)^2 + (y - k)^2 = r^2
Where (h,k) is the center, and 'r' is the radius.
C(2,-3)
radius: 2*sqrt(5)

Answer by jake_6233(74)

Given that (x-h)^2 +(y-c)^2=r^2
(x-2)^2+(y+3)^2=20
So the center is (2,-3) and the radius is the sqrt(20) or 2*sqrt(5)